commercially available h2so4 contains 91% acid by mass and has density of 1.83g/mL calculate--1-molarity of solution and volume of concentrated acid required to prepare 3.5 L of 0.50 M H2SO4

Grade:11

1 Answers

Naveen Kumar
9 years ago
let 100g of acid
So in ths the mass of H2SO4=91g
and volume on which thismuch acid is present=100g/density=(100/1.83)ml=54.64ml
Now the number of moles of acid in this sample=91/MH2SO4=91/98 mol
Hence molarity={Moles/Volume(ml)}*1000=(91/98)*1000/54.64 M
Now if we have the molarity of original solution so apply M1V1=M2V2 to get the required volume of the acd t make desired concentration of acid …...............................